OpenStudy (anonymous):

what is sin (7pi/4) ?

OpenStudy (anonymous):

\[\sin \frac{ 7\pi }{ 4 }\]

OpenStudy (galleara):

since sin \[\pi/4=\sqrt{2}/2\] and 7pi/4 is in quadrant 4, the answer is\[-\sqrt{2}/2\]

OpenStudy (anonymous):

thank u

OpenStudy (galleara):

you're welcome :)